# Algebraic Topology Principles behind Topological Quantum Error Correction

### v1

@@ V Examples of TQEC Codes on 2-Manifolds and Simulation Results
V.1 Implementation of the TQEC Codes
@@
The toric code is defined on a two-dimensional square lattice with periodic boundary conditions, effectively forming a torus. The arrangement of qubits is depicted in Figure 10 [21, 23]. As discussed previously, logical $\sigma_{X}$ operations are implemented by applying physical $\sigma_{X}$ operators along a loop that wraps around the torus. Mathematically, this corresponds to acting with physical $\sigma_{X}$ operators on $[\gamma]\in C_{1}$, where $[\gamma]\neq 0$ and $\partial[\gamma]=0$. This construction highlights the topological nature of the code, as the logical operators are associated with non-trivial homological cycles on the underlying manifold.
@@
Recalled from the examples above, $H_{1}(K;\mathbb{Z}_{2})=\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$, and $H_{1}(\mathbb{R}P^{2};\mathbb{Z}_{2})=\mathbb{Z}_{2}$. Consequently, only a single qubit can be encoded when utilizing the topology of $\mathbb{R}P^{2}$. In contrast, the Klein bottle topology allows encoding two qubits, similar to the toric code. Therefore, in subsequent sections, we will present simulations for the quantum error correction code based on the Klein bottle topology. This analysis evaluates its performance and potential advantages compared to the original toric code.
V.2 Methodology
@@
V.3 Simulation with Random Qubit Errors
@@
Figure 13: Logical versus physical $\sigma_{Z}$ error rates for TQEC code embedded on torus $T^{2}$ and Klein bottle $K$ for different code distances assuming perfect syndrome measurements. (a) The physical error rate range from $1\%-20\%$ for $d=\{10,20,40,80\}$. (b) The physical error rate range from $3\%-10\%$ for $d=\{10,20,40,80\}$. (c) The physical error rate range from $3\%-10\%$ for $d=\{4,6,8\}$. The TQEC code for qubits on $K$ outperforms that on $T^{2}$ when $d$ is even. Codes with smaller $d$ show more prominent improvements.
@@
Figure 14: Logical versus physical $\sigma_{Z}$ error rates for TQEC code embedded on torus $T^{2}$ and Klein bottle $K$ for different code distances $d=\{5,11,21,41\}$ assuming perfect syndrome measurements. There is no statistically significant distinction between the performance of TQEC codes on $K$ or $T^{2}$ when $d$ is odd.
@@
In the first part of this study, we simulate the physical and logical $\sigma_{Z}$ error rates for qubits on cell complexes embedded in the torus $T^{2}$ and the Klein bottle $K$ using various code distances $d=\{10,20,40,80\}$. The physical error rate and the corresponding logical error rate are collected and plotted. The physical error rate $p$ is sampled from a uniformly distributed random variable. For each code distance $d$, simulations are conducted over 5000 runs for both $T^{2}$ and $K$. The error bars in the plots represent 1-sigma uncertainties, and for simplicity, we assume noiseless syndrome measurements. Simulation results incorporating noisy syndromes will be presented in section V.4.
@@
The outcomes of these simulations are shown in Figures 13(a) and 13(b). These figures depict the relationship between logical and physical error rates for TQEC codes embedded on $T^{2}$ and $K$ across different code distances. Figure 13(a) covers a physical error rate range of $1\%-20\%$, while Figure 13(b) focuses on a range of $3\%-10\%$. The results indicate that TQEC codes on $T^{2}$ and $K$ exhibit similar scaling behavior. However, the logical error rate for codes on the Klein bottle is consistently slightly lower across the considered physical error rates. Interestingly, this improvement is more pronounced for smaller code distances $d$ compared to larger ones, which is even clearer when considering Figure 13(c). A qualitative explanation for these observations is provided in Section V.5.
@@
In all simulations described above, we fix $d$ to even values. To explore the impact of odd code distances, we conducted additional simulations with $d=\{5,11,21,41\}$ over a physical error rate range of $3\%-10\%$. The results are presented in Figure 14. In this case, the TQEC code on the Klein bottle does not demonstrate significant improvement over the Toric code. A qualitative explanation for this behavior is also provided in section V.5.
V.4 Simulation with Random Qubit Errors and Noisy Syndromes
@@
Figure 15: Logical versus physical $\sigma_{Z}$ error rates for TQEC code embedded on torus $T^{2}$ and Klein bottle $K$ for different code distances $d=\{10,20,30\}$ with the phenomenological error model for syndrome measurements. (a) The physical error rate ranges from $2\%-4\%$. (b) The physical error rate ranges from $2\%-3\%$. The TQEC code for qubits on the $K$ outperforms that on $T^{2}$ when $d$ is even. Codes with smaller $d$ show more prominent improvements.
@@
In addition to simulating qubit errors, noisy syndrome measurements are incorporated into the simulations now. These simulations are conducted for various code distances $d=\{10,20,30\}$, with each data point representing the mean of 5000 runs. The error bars in the plots correspond to 1-sigma uncertainties. The simulation methodology involves the modified example codes from [40]. To mitigate errors in syndrome measurements, each measurement is repeated multiple times. Consequently, the decoding process is performed on a three-dimensional lattice instead of a two-dimensional lattice, with the third dimension representing the number of syndrome measurement repetitions [42].
@@
The results are presented in Figures 15(a) and 15(b). A notable observation is that compared to perfect syndrome measurements shown in Figures 13(a) and 13(b), the logical error rate increases significantly faster when syndrome measurement errors are present. Despite this increase, TQEC codes defined on the Klein bottle continue to slightly outperform the toric codes. Furthermore, the improvement is more pronounced for smaller code distances $d$. A qualitative explanation of these observations is provided next.
V.5 Discussion
@@
From the above simulations, it is evident that utilizing Klein bottles for TQEC provides a slight improvement over the original toric code when the lattice dimension is even, both with and without noisy syndromes. Notably, this improvement is more pronounced for smaller lattice sizes or smaller code distances $d$. This observation is particularly intriguing, and in this subsection, we aim to provide a qualitative explanation for this phenomenon.
@@
For this reason, we expect that the improvement offered by using Klein bottles becomes more significant at smaller code distances $d$. Consider a square lattice structure as shown in Figure 10. Sorting all logical $\sigma_{Z}$ operations in ascending order of length yields a sequence $\Gamma=\{\gamma_{1},\gamma_{2},\dots\}$ such that $\gamma_{i}\leq\gamma_{j}$ for $i<j$. For a torus, all logical loops have equal lengths: $\gamma_{1}=\gamma_{2}=\cdots=\gamma_{d}=d$. However, for Klein bottles with code distance $d$, the shortest loop satisfies $\gamma_{1}=d+1$. This idea is illustrated in Figure 16(a), where the shortest homology loop $\beta_{1}\in C^{1}$ is one qubit longer than its counterpart in the toric code. Consequently, we observe more prominent improvements for smaller values of $d$.
@@
An additional noteworthy phenomenon emerges when we consider the implementation presented in Figure 16, which we considered a lattice-structured cell complex embedded on both the torus and the Klein bottle. As illustrated in Figures 16(a) and 16(b), the behavior of the code depends critically on whether the lattice dimension $L$ is even or odd. Specifically, when $L$ is even, the shortest $\gamma\in C_{1}$, represented by the red line in the figures, increases in length by 1 compared to its counterpart in the toric code with the same $L$. This results in an increment of one in the length of the shortest logical $\sigma_{X}$ operator, which is responsible for decoding logical $\sigma_{Z}$ errors. Consequently, as shown in Figures 13(a) and 13(b), the logical $\sigma_{Z}$ error rates exhibit improvement over the toric code when $L$ is even. Conversely, for odd values of $L$, the shortest $\gamma\in C_{1}$ remains identical for both the Klein bottle and the torus, and thus no improvement in logical error rates is observed for odd code distances $d$.
@@
Furthermore, as illustrated in Figures 16(a) and 16(b), the shortest 1-cocycle $\beta\in C^{1}$, depicted as the blue line, has the same length for both the torus $T^{2}$ and Klein bottle $K$ topologies. Since $\beta$ corresponds to the logical $\sigma_{Z}$ operators, which are responsible for detecting logical $\sigma_{X}$ errors, we do not anticipate any improvement in $\sigma_{X}$ error rates when employing the Klein bottle topology, irrespective of whether $L$ is even or odd. The corresponding simulation results are provided in Appendix C.
VI Conclusion
@@
In this paper, we have explored the theoretical foundations of Topological Quantum Error Correction (TQEC) through an algebraic topology approach. We began by reviewing key properties and long-established theorems in algebraic topology, subsequently demonstrating the relationship between TQEC capabilities and intersection properties on manifolds, supported by rigorous proofs. Focusing on 2-manifolds $M$, we established a connection between homology and cohomology groups and intersection properties, deriving the necessary and sufficient conditions for a 2-manifold to serve as a TQEC code. Specifically, we proved that a 2-manifold can be used for TQEC for qubits if and only if $H_{1}(M;\mathbb{Z}_{2})$ is non-trivial, as formalized in Theorem III.3. This theoretical framework was then generalized to higher-dimensional manifolds, which is formalized in Theorem IV.1 stating that an $n$-manifold $M$ can encode qubits on the $i$-th dimension if $H_{i}(M;\mathbb{Z}_{2})$ is non-trivial. Then, we provide a realistic example of defining TQEC codes on the 3-torus with detailed implementation schemes, with performance analysis, illustrating the potential of such higher-dimensional manifolds for TQEC applications. This opens up a much wider possibilities for TQEC codes for qubits.
@@ Acknowledgements.
Xiang Zou would like to express his sincere gratitude to Professor Dror Bar-Natan of the Department of Mathematics, University of Toronto, for his valuable comments and insightful discussions, which have greatly contributed to the development of this work. Xiang Zou is supported through the Scholarships for Doctoral Study by the Croucher Foundation. Hoi-Kwong Lo is supported by MITACS, NSERC, Canadian Foundation for Innovation (CFI), the High Throughput and Secure Networks Challenge Program at the National Research Council of Canada (NRC), and the NRC Applied Quantum Computing Challenge Program.
Appendix A Mathematical Foundation

### v2

@@ VII Examples of TQEC Codes on 2-Manifolds and Simulation Results
VII.1 Implementation of the TQEC Codes
@@
The toric code is defined on a two-dimensional square lattice with periodic boundary conditions, effectively forming a torus. The arrangement of qubits is depicted in Figure 1 [21, 23]. As discussed previously, logical $\sigma_{X}$ operations are implemented by applying physical $\sigma_{X}$ operators along a loop that wraps around the torus. Mathematically, this corresponds to acting with physical $\sigma_{X}$ operators on $\gamma]\in C_{1}$, where $\gamma]\neq 0$ and $\partial[\gamma]=0$. This construction highlights the topological nature of the code, as the logical operators are associated with non-trivial homological cycles on the underlying manifold.
@@
Recalled from the examples above, $H_{1}(K;\mathbb{Z}_{2})=\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$, and $H_{1}(\mathbb{R}P^{2};\mathbb{Z}_{2})=\mathbb{Z}_{2}$. Consequently, only a single qubit can be encoded when utilizing the topology of $\mathbb{R}P^{2}$. In contrast, the Klein bottle topology allows for the encoding of two qubits, similar to the toric code. Therefore, in subsequent sections, we will present simulations for the quantum error correction code based on the Klein bottle topology. This analysis evaluates its performance and potential advantages compared to the original toric code.
VII.2 Methodology
@@
VII.3 Simulation with Random Qubit Errors
@@
Figure 24: Logical versus physical $\sigma_{Z}$ error rates for TQEC code embedded on torus $T^{2}$ and Klein bottle $K$ for different lattice dimensions assuming perfect syndrome measurements. (a) The physical error rate range from $1\%-20\%$ for $L=\{10,20,40,80\}$. (b) The physical error rate range from $3\%-10\%$ for $L=\{10,20,40,80\}$. (c) The physical error rate range from $3\%-10\%$ for $L=\{4,6,8\}$. The TQEC code for qubits on $K$ outperforms that on $T^{2}$ when $d$ is even. Codes with smaller $d$ show more prominent improvements.
@@
Figure 25: Logical versus physical $\sigma_{Z}$ error rates for TQEC code embedded on torus $T^{2}$ and Klein bottle $K$ for different lattice dimensions $L=\{5,11,21,41\}$ assuming perfect syndrome measurements. There is no statistically significant distinction between the performance of TQEC codes on $K$ or $T^{2}$ when $d$ is odd.
@@
In the first part of this study, we simulate the physical and logical $\sigma_{Z}$ error rates for qubits on cell complexes embedded in the torus $T^{2}$ and the Klein bottle $K$ using various lattice dimensions $L=\{10,20,40,80\}$. The physical error rate and the corresponding logical error rate are collected and plotted. The physical error rate $p$ is sampled from a uniformly distributed random variable. For each lattice dimension $L$, simulations are conducted over 5000 runs for both $T^{2}$ and $K$. The error bars in the plots represent 1-sigma uncertainties, and for simplicity, we assume noiseless syndrome measurements. Simulation results incorporating noisy syndromes will be presented in section VII.4.
@@
The outcomes of these simulations are shown in Figures 24(a) and 24(b). These figures depict the relationship between logical and physical error rates for TQEC codes embedded on $T^{2}$ and $K$ across different lattice dimensions. Figure 24(a) covers a physical error rate range of $1\%-20\%$, while Figure 24(b) focuses on a range of $3\%-10\%$. The results indicate that TQEC codes on $T^{2}$ and $K$ exhibit similar scaling behavior. However, the logical error rate for codes on the Klein bottle is consistently slightly lower across the considered physical error rates. Interestingly, this improvement is more pronounced for smaller lattice dimensions $L$ compared to larger ones, which is even clearer when considering Figure 24(c). A qualitative explanation for these observations is provided in Section VII.5.
@@
In all simulations described above, we fix $d$ to even values. To explore the impact of odd lattice dimensions, we conducted additional simulations with $L=\{5,11,21,41\}$ over a physical error rate range of $3\%-10\%$. The results are presented in Figure 25. In this case, the TQEC code on the Klein bottle does not demonstrate significant improvement over the Toric code. A qualitative explanation for this behavior is also provided in section VII.5.
VII.4 Simulation with Random Qubit Errors and Noisy Syndromes
@@
Figure 26: Logical versus physical $\sigma_{Z}$ error rates for TQEC code embedded on torus $T^{2}$ and Klein bottle $K$ for different lattice dimensions $L=\{10,20,30\}$ with the phenomenological error model for syndrome measurements. (a) The physical error rate ranges from $2\%-4\%$. (b) The physical error rate ranges from $2\%-3\%$. The TQEC code for qubits on the $K$ outperforms that on $T^{2}$ when $L$ is even. Codes with smaller $L$ show more prominent improvements.
@@
In addition to simulating qubit errors, noisy syndrome measurements are incorporated into the simulations now. These simulations are conducted for various lattice dimensions $L=\{10,20,30\}$, with each data point representing the mean of 5000 runs. The error bars in the plots correspond to 1-sigma uncertainties. The simulation methodology involves the modified example codes from [56]. To mitigate errors in syndrome measurements, each measurement is repeated multiple times. Consequently, the decoding process is performed on a three-dimensional lattice instead of a two-dimensional lattice, with the third dimension representing the number of syndrome measurement repetitions [58].
@@
The results are presented in Figures 26(a) and 26(b). A notable observation is that compared to perfect syndrome measurements shown in Figures 24(a) and 24(b), the logical error rate increases significantly faster when syndrome measurement errors are present. Despite this increase, TQEC codes defined on the Klein bottle continue to slightly outperform the toric codes. Furthermore, the improvement is more pronounced for smaller lattice dimensions $L$. A qualitative explanation of these observations is provided next.
VII.5 Discussion
@@
From the above simulations, it is evident that utilizing Klein bottles for TQEC provides a slight improvement over the original toric code when the lattice dimension is even, both with and without noisy syndromes. Notably, this improvement is more pronounced for smaller lattice sizes or smaller lattice dimensions $L$. This observation is particularly intriguing, and in this subsection, we aim to provide a qualitative explanation for this phenomenon.
@@
For this reason, we expect that the improvement offered by using Klein bottles becomes more significant at smaller lattice dimensions $L$. Consider a square lattice structure as shown in Figure 1. Sorting all logical $\sigma_{Z}$ operations in ascending order of length yields a sequence $\Gamma=\{\gamma_{1},\gamma_{2},\dots\}$ such that $\gamma_{i}\leq\gamma_{j}$ for $i<j$. For a torus, all logical loops have equal lengths: $\gamma_{1}=\gamma_{2}=\cdots=\gamma_{d}=d$. However, for Klein bottles with lattice dimension $L$, the shortest loop satisfies $\gamma_{1}=d+1$. This idea is illustrated in Figure 27(a), where the shortest homology loop $\beta_{1}\in C^{1}$ is one qubit longer than its counterpart in the toric code. Consequently, we observe more prominent improvements for smaller values of $d$.
@@
An additional noteworthy phenomenon emerges when we consider the implementation presented in Figure 27, which we considered a lattice-structured cell complex embedded on both the torus and the Klein bottle. As illustrated in Figures 27(a) and 27(b), the behavior of the code depends critically on whether the lattice dimension $L$ is even or odd. Specifically, when $L$ is even, the shortest $\gamma\in C_{1}$, represented by the red line in the figures, increases in length by 1 compared to its counterpart in the toric code with the same $L$. This results in an increment of one in the length of the shortest logical $\sigma_{X}$ operator, which is responsible for decoding logical $\sigma_{Z}$ errors. Consequently, as shown in Figures 24(a) and 24(b), the logical $\sigma_{Z}$ error rates exhibit improvement over the toric code when $L$ is even. Conversely, for odd values of $L$, the shortest $\gamma\in C_{1}$ remains identical for both the Klein bottle and the torus, and thus no improvement in logical error rates is observed for odd lattice dimensions $L$.
@@
Furthermore, as illustrated in Figures 27(a) and 27(b), the shortest 1-cocycle $\beta\in C^{1}$, depicted as the blue line, has the same length for both the torus $T^{2}$ and Klein bottle $K$ topologies. Since $\beta$ corresponds to the logical $\sigma_{Z}$ operators, which are responsible for detecting logical $\sigma_{X}$ errors, we do not anticipate any improvement in $\sigma_{X}$ error rates when employing the Klein bottle topology, irrespective of whether $L$ is even or odd. The corresponding simulation results are provided in Appendix C.
VIII Conclusion
@@
In this paper, we have explored the theoretical foundations of Topological Quantum Error Correction (TQEC) through an algebraic topology approach. We began by reviewing key properties and long-established theorems in algebraic topology, subsequently demonstrating the relationship between TQEC capabilities and intersection properties on manifolds, supported by rigorous proofs. Focusing on 2-manifolds $M$, we established a connection between homology and cohomology groups and intersection properties, deriving the necessary and sufficient conditions for a 2-manifold to serve as a TQEC code. Specifically, we proved that a 2-manifold can be used for TQEC for qubits if and only if $H_{1}(M;\mathbb{Z}_{2})$ is non-trivial, as formalized in Theorem III.3. This theoretical framework was then generalized to higher-dimensional manifolds, which is formalized in Theorem IV.1 stating that an $n$-manifold $M$ can encode qubits on the $i$-th dimension if $H_{i}(M;\mathbb{Z}_{2})$ is non-trivial. Then, we provide a realistic example of defining TQEC codes on the 3-torus with detailed implementation schemes, with performance analysis, illustrating the potential of such higher-dimensional manifolds for TQEC applications. This opens up many wider possibilities for TQEC codes for qubits.
@@ Acknowledgements.
Xiang Zou expresses his sincere gratitude to Professor Dror Bar-Natan of the Department of Mathematics, University of Toronto, for his valuable comments and insightful discussions, which have greatly contributed to the development of this work. Xiang Zou is supported through the Scholarships for Doctoral Study by the Croucher Foundation. Hoi-Kwong Lo is supported by MITACS, NSERC, Canadian Foundation for Innovation (CFI), the High Throughput and Secure Networks Challenge Program at the National Research Council of Canada (NRC), the NRC Applied Quantum Computing Challenge Program, NUS start-up grant, and CQT Core Research Funding.
IX Data Availability
The code used for the simulations and all presented data are available upon request.
X Author Contributions
Xiang Zou contributed to the theorem, proofs, and examples in the article. Hoi-Kwong Lo conceptualized, proofread, and modified the work.
Appendix A Proofs of Theorems
For readability, we have omitted most of the proofs of the theorems, and in this section, the omitted proofs are provided.
Theorem III.2.
Firstly, using the basic principle of QEC, by proper stabilizer measurements can collapse the set of all possible quantum errors t of discrete quantum errors. Thus, for any possible qubit error $Err$, it can be written as $Err=e_{0}I+e_{1}\sigma_{X}+e_{2}\sigma_{Z}+e_{3}\sigma_{X}\sigma_{Z}$ for some constant $e_{0},e_{1},e_{2},e_{3}$. This means, for QEC purposes, a method to implement logical $\sigma_{X}$ and $\sigma_{Z}$ on the code is sufficient.
Corollary III.2.1.
Theorem III.2 states that, for any closed surfaces to be used as a TQEC code, we require at least 2 loops that interact in an odd number of points. Theorem III.1 states that the modulo-2 intersection number is homotopic invariant. However, for any 2 non-overlapping loops on simply connected surfaces, each can be shrunk down to a point continuously, where they do not intersect. This means that two non-overlapping loops can only transversely intersect at an even number of points. Thus, they cannot be used as TQEC codes. ∎
Theorem III.3.
Notation-wise, since $\mathbb{Z}_{2}$ is commutative, we will represent it as the modulo-2 addition group that the elements of $\mathbb{Z}_{2}$ are $\{0,1\}$, and for simplicity, for $a\in\{0,1\}$ we represent $\sum_{i=1}^{n}a:=a^{n}$.
Then, to prove the “if” part, considering elements belonging to the $H_{1}(M;\mathbb{Z}_{2})$, $\gamma\in C_{1}(M;\mathbb{Z}_{2}),[\gamma]\neq 0\text{ and }\partial\gamma=0$. And elements belongs to the $H^{1}(M;\mathbb{Z}_{2})$, $\beta\in C^{1}(M;\mathbb{Z}_{2}),[\beta]\neq 0\text{ and }\delta\beta=0$. The Poincaré duality as in the theorem II.1 states that $H_{1}(M;\mathbb{Z}_{2})\cong H^{1}(M;\mathbb{Z}_{2})$. Thus, the pairing given by a function $p:H_{1}(M;\mathbb{Z}_{2})\times H^{1}(M;\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}$ is well-defined, and the $\mathbb{Z}_{2}$ coefficients ensure that $\gamma]^{2}=0$ and $\beta]^{2}=0$. By the properties of elements belonging to the first cohomology group, for arbitrary integers $r,s$, we have $\beta]^{2r+1}([\gamma]^{2s+1})=[\beta]([\gamma])=1$, while $\beta]^{2r}([\gamma]^{s})=[\beta]^{r}([\gamma]^{2s})=0$. The Poincaré duality states that since $H_{1}(M;\mathbb{Z}_{2})$ is isomorphic to $H^{1}(M;\mathbb{Z}_{2})$, thus, there is a bijective mapping between elements in $C^{1}$ to elements in $C_{1}$, so since $\gamma]$ is a class of closed loops, $\beta]$ are also closed loops on the surface $M$. Therefore, logical $\sigma_{X}$ can be defined on $\gamma]$ and logical $\sigma_{Z}$ can be defined on $\beta]$, where they anti-commute, which defines a TQEC code.
Theorem IV.1.
Firstly, we prove that the closed and compact $n$-manifold $M$ can be used for a TQEC code for qubits on the $i$-cycle and $i$-cocycle only if its $i$-th homology group is nontrivial. The reason being that if $H_{i}(M;\mathbb{Z}_{2})=0$, then $b_{i}(M;\mathbb{Z}_{2})=0$. There is only 1 element in the group $H_{i}(M;\mathbb{Z}_{2})$, which means all closed $i$-cycles are equivalently null-homologous. This implies all $\gamma\in C_{i}(M;\mathbb{Z}_{2}),\partial\gamma=0$ are equivalent, that $\gamma^{s}=\gamma$ for all positive integer $s$. By the definition of elements belonging to the cohomology group $H^{i}(M;\mathbb{Z}_{2})$, even if there exist $\beta\in C^{i}(M;\mathbb{Z}_{2}),[\beta]\neq 0\text{ and }\delta\beta=0$. For all non-negative integer $r$, $\beta(\gamma^{r})=\beta(\gamma)=0$. However, for $\sigma_{X}$ operator, $\sigma_{X}^{2r+1}=\sigma_{X}$ and $\sigma_{X}^{2r}=\mathbb{I}$ for an arbitrary integer $r$. This implies that $\sigma_{X}$ cannot be defined on any $i$-cycles; hence, defining a TQEC code is impossible.
Then, to prove the “if” part, considering elements belonging to the $H_{i}(M;\mathbb{Z}_{2})$, $\gamma\in C_{i}(M;\mathbb{Z}_{2}),[\gamma]\neq 0\text{ and }\partial\gamma=0$. And elements belongs to the $H^{i}(M;\mathbb{Z}_{2})$, $\beta\in C^{i}(M;\mathbb{Z}_{2}),[\beta]\neq 0\text{ and }\delta\beta=0$. The Poincaré duality as in the theorem II.1 states that $H_{i}(M;\mathbb{Z}_{2})\cong H^{n-i}(M;\mathbb{Z}_{2})$. Thus, the pairing given by a function $p:H_{i}(M;\mathbb{Z}_{2})\times H^{i}(M;\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}$ is well-defined. The $\mathbb{Z}_{2}$ coefficients ensure that $\gamma]^{2}=[\beta]^{2}=0$. And by the definition of elements belonging to the cohomology group, for arbitrary integers $r,s$, we have $\beta]^{2r+1}([\gamma]^{2s+1})=[\beta]([\gamma])=1$, while $\beta]^{2r}([\gamma]^{s})=[\beta]^{r}([\gamma]^{2s})=0$. $\gamma]$ represent the class of non-trivial closed $i$-cycles and $\beta]$ represent the class of non-trivial closed $i$-cocycles, which is isomorphic to the class of non-trivial closed (ni)(n-i)-cycles. Therefore, logical $\sigma_{X}$ can be defined on $\gamma]$ and logical $\sigma_{Z}$ can be defined on $\beta]$, where they anti-commute, which defines a TQEC code.
Theorem V.1.
The proof is similar to that of Theorem IV.1. Given a compact $n$-manifold $M$ with boundary $\partial M=A\cup B$ and $\partial A=\partial B=A\cap B$, the reason why we decompose $\partial M$ into two parts $A$ and $B=\partial M\setminus A\cup\partial A$ is that, for qubit codes, there are only 2 types of boundary conditions applied, $x$-type boundary and $z$-type boundary as proven by Kitaev in Ref.[22, 45]. Thus, when we have $\partial M=A\cup B$, we may use $A$ as the $x$-boundary and $B$ as the $z$-boundary or vice versa, then the proof follows.
Firstly, we prove that $M$ can be used for a TQEC code for qubits on the $i$-cycle and $i$-cocycle only if $\exists A,s.t.,H_{i}(M,A;\mathbb{Z}_{2})\neq 0$. The reason being that if $\forall A,H_{i}(M,A;\mathbb{Z}_{2})=0$, then $\forall A,b_{i}(M,A;\mathbb{Z}_{2})=0$. There is only 1 element in the group $H_{i}(M,A;\mathbb{Z}_{2}),\forall A$. Which means all closed $i$-cycles are equivalently null-homologous. This implies all $\gamma\in C_{i}(M,A;\mathbb{Z}_{2}),\partial\gamma=0,\forall A$ are equivalent, that $\gamma^{s}=\gamma$ for all positive integer $s$. This implies that $\sigma_{X}$ cannot be defined on any $i$-cycles; hence, defining a TQEC code is impossible.
Then, to prove the “if” part, suppose $\exists A,s.t.,H_{i}(M,A;\mathbb{Z}_{2})\neq 0\text{ and }H_{n-i}(M,B;\mathbb{Z}_{2})$. Considering elements belonging to the $H_{i}(M,A;\mathbb{Z}_{2})$, $\gamma\in C_{i}(M,A;\mathbb{Z}_{2}),[\gamma]\neq 0\text{ and }\partial\gamma=0$. And elements belongs to the $H^{i}(M,A;\mathbb{Z}_{2})$, $\beta\in C^{i}(M;\mathbb{Z}_{2}),[\beta]\neq 0\text{ and }\delta\beta=0$. The Lefschetz duality in Theorem II.2.1 states that $H^{i}(M,A;\mathbb{Z}_{2})\cong H_{n-i}(M,B;\mathbb{Z}_{2})$. Thus, the pairing given by a function $p:H_{i}(M,A;\mathbb{Z}_{2})\times H^{i}(M,A;\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}$ is well-defined. The $\mathbb{Z}_{2}$ coefficients ensure that $\gamma]^{2}=[\beta]^{2}=0$. And by the definition of elements belonging to the cohomology group, for arbitrary integers $r,s$, we have $\beta]^{2r+1}([\gamma]^{2s+1})=[\beta]([\gamma])=1$, while $\beta]^{2r}([\gamma]^{s})=[\beta]^{r}([\gamma]^{2s})=0$. $\gamma]$ represent the class of non-trivial closed $i$-cycles and $\beta]$ represents the class of non-trivial closed $i$-cocycles with the presence of boundary $B$, which is isomorphic to the class of non-trivial closed (ni)(n-i)-cycles with the presence of boundary $A$. Therefore, logical $\sigma_{X}$ can be defined on $\gamma]$ and logical $\sigma_{Z}$ can be defined on $\beta]$, where they anti-commute, which defines a TQEC code.
Since from the Lefschetz duality in Theorem II.2.1, $H^{i}(M,A;\mathbb{Z}_{2})\cong H_{n-i}(M,B;\mathbb{Z}_{2})$. The rank of $H_{i}(M,A;\mathbb{Z}_{2})$ determines the number of different classes of $i$-chains that can be defined on $M$, and the number of different classes of $i$-cochain equals the rank of $H_{n-i}(M,B;\mathbb{Z}_{2})$. Which means, the number of qubits encoded is determined by the smaller of $b_{i}(M,A;\mathbb{Z}_{2})$ and $b_{n-i}(M,B;\mathbb{Z}_{2})$. ∎
Theorem VI.2.
We will prove that such a presentation complex $X_{G}=(V,E,F)$ built from a group $G=\mathbb{Z}_{d}^{k}$ can encode a $d$-dimensional qudits by providing a way to define the logical $X^{a}$ and $Z^{b}$ operators of the qudits.
From the construction of a presentation complex from a group $\mathbb{Z}_{d}$, the idea is to use a d-pointed “asterisk” with a $1/d$ twist after a revolution, just as shown in figure 17, and then attach a disk along the boundary circle of the d-pointed “asterisk”. The logical $X$ and logical $Z$ are defined to be closed loops on the presentation complex; they wrap around the presentation complex in two different directions, as shown in Figure 20.
Similar to the normal qubit code, the Hamiltonian is defined to be the summation of all stabilizers defined in Figure 19, and both the logical $X$ and $Z$ operators commute with the Hamiltonian 33. Just a reminder, the single-qudit operator $X$ and $Z$ are defined in equation 29. So, $X^{a}Z^{b}=\omega^{ab}Z^{b}X^{a}$, for $\omega=e^{2\pi i/d}$.
In other words, the logical $X$ operates on the homological cycles $C_{1}(X_{G})$, $\gamma\in C_{1}(X_{G};\mathbb{Z}_{d}),[\gamma]\neq 0\text{ and }\partial\gamma=0$, while the logical $Z$ operates on the co-cycles $C^{1}(X_{G})$, $\beta\in C^{1}(X_{G};\mathbb{Z}_{d}),[\beta]\neq 0\text{ and }\delta\beta=0$. Since $H_{1}(X_{G})=\mathbb{Z}_{d}$, there should be $d$ distinct elements in $C_{1}(X_{G};\mathbb{Z}_{d})$, which corresponds to revolving around the presentation complex $0,\cdots,d-1$ times. Hence, the logical $X^{a}$ operates on a $\gamma\in C_{1}(X_{G};\mathbb{Z}_{d}),[\gamma]=(a\mod d)$, so for any logical $X^{a}$, there exist a $\gamma\in C_{1}(X_{G};\mathbb{Z}_{d})$.
By the properties of the presentation complex and group elements belonging to the first homology and cohomology groups [26], for an arbitrary power of the operator $X^{r}$ and $Z^{s}$, denoting $a=r\mod d,b=s\mod d$. We have $Z]^{r}([X]^{s})=[Z]^{a}([X])^{b}$, where the two loops intersects at $ab$ number of sites. Thus, the exchange statistics will be $X^{a}Z^{b}=\omega^{ab}Z^{b}X^{a}$.
[…]
Hence,
[…]
Hence, the above definition of the operators works as the logical $X$ and logical $Z$ for the encoded qudit.
Now, the only problem left is whether, when the logical operators get detoured, any of the above arguments will be violated? The answer is no. Remember that the sign ($\pm 1$) of each intersection point is well-defined since the presentation complex is oriented. For two non-detoured logical operators $X$ and $Z$ intersecting at 1 site, we define this intersection point to have a sign +1+1.
Up to any curve homotopy, the number of intersection points will only be increased by an even number; this is a well-known result [26, 27], and among the newly emerged intersection points, always half of them have a sign of +1+1 and half are 1-1, this is also a well-known textbook result [26, 27]. This means, in the detoured case, if 2 intersection points are added, one of them would be an intersection between $\{X,Z^{-1}\}$ or $\{X^{-1},Z\}$, while the other intersection would be between $\{X,Z\}$ or $\{X^{-1},Z^{-1}\}$, then the two intersection will “cancel out”, and not affecting the overall phase of the logical operators.
Appendix B Mathematical Foundation